OpenStudy (anonymous):

Find the solution of y"+2y'+2y=delta(t-pi); y(0)=1, y'(0)=0.

OpenStudy (anonymous):

delta(t-π) is the dirac delta function, correct?

OpenStudy (anonymous):

Take the Laplace transform of both sides: \[\left(s^2Y(s)-sy(0)-y'(0)\right)+2\left(sY(s)-y(0)\right)+2Y(s)=e^{-\pi s}\\ \left(s^2Y(s)+2sY(s)+2Y(s)\right)-s-2=e^{-\pi s}\\ Y(s)\left(s^2+2s+2\right)=e^{-\pi s}+s+2\\ Y(s)=\frac{e^{-\pi s}+s+2}{s^2+2s+2}\\ Y(s)=\frac{e^{-\pi s}}{s^2+2s+2}+\frac{s}{s^2+2s+2}+\frac{2}{s^2+2s+2}\] It should be easy to find the inverse transforms.

OpenStudy (anonymous):

Complete the square in each denominator: \[s^2+2s+2=(s+1)^2+1\] It might be easier to first write \[Y(s)=\frac{e^{-\pi s}}{s^2+2s+2}+\frac{s+1}{s^2+2s+2}+\frac{1}{s^2+2s+2}\] So you have \[Y(s)=\frac{e^{-\pi s}}{(s+1)^2+1}+\frac{s+1}{(s+1)^2+1}+\frac{1}{(s+1)^2+1}\]

OpenStudy (anonymous):

Thank you.